OpenStudy (anonymous):

Carrie leaves her home at 5:10 p.m and drives to her sisters house . She drives 16 miles and arrives at her sisters house at 5:34 p.m. What is carries average speed in miles per minute

OpenStudy (anonymous):

well it takes 24 mins to get to her sisters house

OpenStudy (anonymous):

5:34-5:10 = 24minute = 2/3 hours -elapsed time S= v*t v = S/t v = 16 * 3/2= 24 mph

OpenStudy (anonymous):

I'm sory 5:34-5:10 = 24minute = 2/5 hours S = 16 *5/2 = 40 mph